OpenStudy (anonymous):

What is the slope between the two points (8, 15) and (3, 5)?

OpenStudy (chriss):

\[\frac{5-15}{3-8}\] \[\frac{-10}{-5}\] slope=2

OpenStudy (anonymous):

use the slope formula: Y2-Y1/X2-X1. So (5-15)/(3-8)= -10/-5=2 (when you divide two negitive numbers, the answer will be positive.)